OpenStudy (anonymous):

What is the t-distribution if u have a 95% confidence level and 84 degrees of freedom?

OpenStudy (anonymous):

95% confidence interval n=85-1 = 84 \[t _{a/2}=t _{?/2}\] \[t _{?/2}=t _{?}\] \[t _{?}=?\]